Find the the line of intersection of the two planes 3x - 2y + z=1 2x+4-32=3
Answer #1

Given 3x-2y+z=i o 23t9-32=3 -7 li - 근 = 1+2y-37 in li 2xty - 3+64 +1=3 11a*7y=6 112=6654 - x=6+59 Il 645 inz z = Itzy 18+15